# LatticeVision: Image to Image Networks for Modeling Non-Stationary Spatial Data

### v1

@@ 9 Conclusion
We introduce LatticeVision, a global, image-to-image (I2I) framework for the estimation and emulation of non-stationary spatial processes. By representing both the spatial fields and parameters as images, we employ vision networks to simultaneously estimate all parameters in a single forward pass. We develop a novel, non-stationary data generation pipeline for training such networks without relying on a large, existing corpus of application-specific data. As compared to local approaches, I2I networks demonstrate improvements in accuracy, robustness with few replicates, estimation speed, and ability to capture long-range, anisotropic correlations. We pair these estimators with the LatticeKrig SAR model, enabling fast simulation of large ensembles for non-stationary spatial data.
@@ A.3 Anisotropic extension
To incorporate geometric anisotropy we replace the Laplacian in (8) by the generalised Laplacian $\nabla\cdot H\nabla$, where the positive–definite dispersion matrix $H\in\mathbb{R}^{2\times 2}$ is
$H\;=\;R^{\top}D^{2}R,\qquad R=\begin{bmatrix}\cos\theta&-\sin\theta\\[2.0pt] \sin\theta&\cos\theta\end{bmatrix},\quad D=\begin{bmatrix}\sqrt{\rho}&0\\[2.0pt] 0&1/\sqrt{\rho}\end{bmatrix}.$ (12)
Here, $R$ is a rotation matrix parametrized by $\theta$ and $D$ is a scaling matrix parametrized by $\rho$. For a constant $H$ the anisotropic SPDE becomes
$\bigl{(}\kappa^{2}-\nabla\cdot H\nabla\bigr{)}\,f(\mathbf{s})=\mathcal{W}(\mathbf{s}).$ (13)
Using unit grid spacing and second-order central differences for the second–derivative terms,
$\nabla\cdot H\nabla f_{i,j}\;\approx\;H_{1,1}(f_{i+1,j}-2f_{i,j}+f_{i-1,j})+H_{2,2}(f_{i,j+1}-2f_{i,j}+f_{i,j-1})\\ +\frac{H_{1,2}}{2}\!\bigl{[}f_{i+1,j+1}-f_{i+1,j-1}-f_{i-1,j+1}+f_{i-1,j-1}\bigr{]}.$ (14)
@@
$\Bigl{(}\kappa^{2}+2H_{1,1}+2H_{2,2}\Bigr{)}f_{i,j}\;-\;H_{1,1}(f_{i+1,j}+f_{i-1,j})\;-\;H_{2,2}(f_{i,j+1}+f_{i,j-1})\\ \;-\;\frac{H_{1,2}}{2}\bigl{[}f_{i+1,j+1}-f_{i+1,j-1}-f_{i-1,j+1}+f_{i-1,j-1}\bigr{]}\;=\;e_{i,j}.$ (15)
@@
$\begin{array}[]{c|c|c}\dfrac{H_{1,2}}{2}&-\,H_{2,2}&-\,\dfrac{H_{1,2}}{2}\\ \hline\cr-\,H_{1,1}&\kappa^{2}+2H_{1,1}+2H_{2,2}&-\,H_{1,1}\\ \hline\cr-\,\dfrac{H_{1,2}}{2}&-\,H_{2,2}&\dfrac{H_{1,2}}{2}.\end{array}$ (16)
One incorporates non-stationarity by allowing $H$ and $\kappa^{2}$ to vary in space, as shown in Equation (3).
@@ B.1 Spatial patterns
Coastline: A sharp sigmoidal boundary, perturbed by sinusoidal "bumps".
@@
  • Hyperparameters: $\boldsymbol{\Omega}=(a_{1},a_{2},\lambda_{1},\lambda_{2},x_{1},y_{1},x_{2},y_{2})$, with amplitudes, widths, and locations sampled as previously described in the above ‘‘Bump’’ configuration, and with centers $(x_{1},y_{1}),(x_{2},y_{2})$ being randomly sampled locations within the spatial domain.
@@
  • Hyperparameters: $w_{1}\sim\mathcal{U}(0.1,0.9),w_{2}\sim\mathcal{U}(0.1,1-w_{1})$, and each coastline has its own, independent $(\alpha,\beta,\omega,\gamma)$ parameters as described in the above ‘‘Coastline’’ configuration.
@@
    $p(\mathbf{s})=\begin{cases}p_{\text{min}},p_{\text{max}}\text{ rescaled field},&\text{(min-max scaling)}\\ p_{\text{constant}}\times(1\pm g(\mathbf{s})),&\text{(perturbation scaling)}\end{cases}$ (24)
@@ B.3 Resulting synthetic data
We repeat the process of selecting a pattern, sampling the pattern hyperparameters $\boldsymbol{\Omega}$, sampling the parameter values, and generating the resulting parameter field for $\kappa^{2}(\mathbf{s}),\rho(\mathbf{s}),$ and $\theta(\mathbf{s})$. We then encode all parameter fields into the SAR matrix $B$, draw white noise $\mathbf{e}\sim\mathcal{N}(0,I)$ and solve $B\,\mathbf{y}=\mathbf{e}$. We repeat this for $M$ independent draws of the white noise, resulting in a small ensemble $Y=\{\mathbf{y}^{(m)}\}^{M}_{m=1}$, with each field having the same covariance structure. We process these synthetic fields identically to the way we process the input ESM fields: we perform pixelwise standardization to ensure each pixel has a mean of 0 and a standard deviation of 1. The non-stationarity in our problem does not allow for more sophisticated normalization techniques [LKNormalizing], which either become computationally intractable, require a stationary structure, or are not applicable in this setting. An example spatial field and its associated parameter fields can be seen in Figure 6.
@@ C.2 Training
All networks are implemented in PyTorch [paszke2019pytorch] and trained on a single NVIDIA RTX A6000 GPU. The networks are trained using the AdamW optimizer [adamw] for 200 epochs with a step-wise learning rate decay and early stopping after 10 epochs without validation improvement. We use mean squared error (MSE) loss, computed on normalized parameter values within the training loop. This avoids loss imbalance caused by parameter scale differences while requiring less pre or post-processing from the user.
@@ Appendix D Further experiments with simulated data
RMSE $\downarrow$ MAE $\downarrow$ SSIM $(\uparrow\text{to }1)$ PSNR $\uparrow$ NRMSE $\downarrow$
@@
RMSE $\downarrow$ MAE $\downarrow$ SSIM $(\uparrow\text{to }1)$ PSNR $\uparrow$ NRMSE $\downarrow$
@@
RMSE $\downarrow$ MAE $\downarrow$ SSIM $(\uparrow\text{to }1)$ PSNR $\uparrow$ NRMSE $\downarrow$
@@ Appendix E Further climate application results
We pass the standardized, 30-member ensemble of CESM1 LENS temperature sensitivity anomalies into each of our estimation networks: STUN, UNet, ViT, CNN25, CNN17, and CNN9. Using the predicted parameters, we generate 1,000 synthetic fields from each model. Generation is performed using the LatticeKrig R package [LKrigpackage], which accommodates cylindrical geometry and ensures the correct treatment of spatial distances under the Mercator projection.
To evaluate how well each synthetic ensemble preserves the spatial correlation structure, we randomly select 50 anchor locations and compute the corresponding 50 rows from each ensemble’s empirical correlation matrix. These are then compared to the 50 rows from the original CESM1 ensemble’s correlation matrix using root mean squared error (RMSE). We systematically pair each I2I network with a local CNN of comparable rank (e.g., STUN vs. CNN25, UNet vs. CNN17), and compare their RMSEs across the same anchor locations. As the anchor locations are fixed for all pairings, we conduct paired t-tests between matched RMSEs to compare performance.
Our null hypothesis for each paired comparison is that the mean RMSE difference equals zero, $H_{0}\!:\,\mu_{d}=0$, where $d=\text{RMSE}_{\text{I2I}}-\text{RMSE}_{\text{CNN}}$. The one-sided alternative, $H_{1}\!:\,\mu_{d}<0$, states that the I2I emulator attains a lower RMSE than its CNN counterpart. All tests are conducted with a significance level $\alpha=0.01$. Results are summarized in Table 5. For every comparison, the null hypothesis is rejected and the $99\%$ confidence interval is entirely negative, demonstrating that each I2I network produces significantly more accurate correlation structure estimates than its corresponding local CNN.
@@
Model pair $\mu_{\text{I2I}}$downarrow $\mu_{\text{local}}$downarrow $\mu_{d}$ 99% CI $p_{\text{value}}$
STUN vs. CNN25 0.229 0.243 $-0.013$ ($\infty$- ∞, $-0.007$] $1.1\times 10^{-6}$
UNet vs. CNN17 0.230 0.239 $-0.008$ ($\infty$- ∞, $-0.004$] $3.4\times 10^{-5}$
ViT vs. CNN9 0.228 0.244 $-0.016$ ($\infty$- ∞, $-0.008$] $5.9\times 10^{-6}$
Additionally, we provide analogous visual comparisons to Figure 5 for the UNet, CNN17, ViT, and CNN9 networks in Figure 7 and Figure 8, respectively. These figures highlight a few consistent trends: I2I emulators more accurately preserve the spatial correlation structure, particularly zonal features along the equator, while local CNNs tend to produce over-smoothed correlations with inflated meridional spread.
Figure 7: Top: Standardized temperature fields drawn from the CESM1-LENS ensemble (left), the UNet-based emulator (middle), and the CNN17-based emulator (right). Bottom: Correlations with a chosen location in the Niño 3.4 region at ($212^{\circ}$, $1^{\circ}$) for the same three ensembles.
Figure 8: Top: Standardized temperature fields drawn from the CESM1-LENS ensemble (left), the ViT-based emulator (middle), and the CNN9-based emulator (right). Bottom: Correlations with a chosen location in the Niño 3.4 region at ($212^{\circ}$, $1^{\circ}$) for the same three ensembles.

### v2

@@ 9 CONCLUSION
We introduce LatticeVision, a global, image-to-image (I2I) framework for the estimation and emulation of non-stationary spatial processes. By representing both the spatial fields and parameters as images, we use I2I networks to estimate all parameters simultaneously, in a single forward pass. We also develop a novel pipeline for generating non-stationary training data. We show that I2I networks demonstrate improvements in accuracy, robustness with few replicates, estimation speed, and ability to capture long-range, anisotropic correlations, as compared to local approaches. We pair I2I parameter estimators with the LatticeKrig SAR model, enabling fast simulation of large ensembles for non-stationary spatial data.
@@ A.3 Anisotropic Extension
To incorporate geometric anisotropy we replace the Laplacian in (8) by the generalised Laplacian $\nabla\cdot D\nabla$, where the positive–definite dispersion matrix $D\in\mathbb{R}^{2\times 2}$ is
$D\;=\;R^{\top}\Lambda R,\qquad R=\begin{bmatrix}\cos\theta&-\sin\theta\\[2.0pt] \sin\theta&\cos\theta\end{bmatrix},\quad\Lambda=\begin{bmatrix}\rho&0\\[2.0pt] 0&\frac{1}{\rho}\end{bmatrix}$. (12)
Here, $R$ is a rotation matrix parametrized by $\theta$ and $\Lambda$ is a scaling matrix parametrized by $\rho$. For a constant $D$ the anisotropic SPDE becomes
$\bigl(\kappa^{2}-\nabla\cdot D\nabla\bigr)\,f(\mathbf{s})=\mathcal{W}(\mathbf{s}).$ (13)
Using unit grid spacing and second‑order central differences for the second–derivative terms,
$\nabla\cdot D\nabla f_{i,j}\;\approx\;D_{1,1}(f_{i+1,j}-2f_{i,j}+f_{i-1,j})+D_{2,2}(f_{i,j+1}-2f_{i,j}+f_{i,j-1})\\ +\frac{D_{1,2}}{2}\!\bigl[f_{i+1,j+1}-f_{i+1,j-1}-f_{i-1,j+1}+f_{i-1,j-1}\bigr].$ (14)
@@
$\Bigl(\kappa^{2}+2D_{1,1}+2D_{2,2}\Bigr)f_{i,j}\;-\;D_{1,1}(f_{i+1,j}+f_{i-1,j})\;-\;D_{2,2}(f_{i,j+1}+f_{i,j-1})\\ \;-\;\frac{D_{1,2}}{2}\bigl[f_{i+1,j+1}-f_{i+1,j-1}-f_{i-1,j+1}+f_{i-1,j-1}\bigr]\;=\;e_{i,j}.$ (15)
@@
$\begin{array}[]{c|c|c}\dfrac{D_{1,2}}{2}&-\,D_{2,2}&-\,\dfrac{D_{1,2}}{2}\\ \hline\cr-\,D_{1,1}&\kappa^{2}+2D_{1,1}+2D_{2,2}&-\,D_{1,1}\\ \hline\cr-\,\dfrac{D_{1,2}}{2}&-\,D_{2,2}&\dfrac{D_{1,2}}{2}.\end{array}$ (16)
One incorporates non-stationarity by allowing $D$ and $\kappa^{2}$ to vary in space, as shown in Equation (3).
@@ B.1 Spatial Patterns
Coastline: A sharp sigmoidal boundary, perturbed by sinusoidal ”bumps”.
@@
  • Hyperparameters: $\boldsymbol{\Omega}=(a_{1},a_{2},\lambda_{1},\lambda_{2},x_{1},y_{1},x_{2},y_{2})$, with amplitudes, widths, and locations sampled as previously described in the above “Bump” configuration, and with centers $(x_{1},y_{1}),(x_{2},y_{2})$ being randomly sampled locations within the spatial domain.
@@
  • Hyperparameters: $w_{1}\sim\mathcal{U}(0.1,0.9),w_{2}\sim\mathcal{U}(0.1,1-w_{1})$, and each coastline has its own, independent $(\alpha,\beta,\omega,\gamma)$ parameters as described in the above “Coastline” configuration.
@@
[…]
@@ B.3 Resulting Synthetic Data
We repeat the process of selecting a pattern, sampling the pattern hyperparameters $\boldsymbol{\Omega}$, sampling the parameter values, and generating the resulting parameter field for $\kappa^{2}(\mathbf{s}),\rho(\mathbf{s}),$ and $\theta(\mathbf{s})$. We then encode all parameter fields into the SAR matrix $B$, draw white noise $\mathbf{e}\sim\mathcal{N}(0,I)$ and solve $B\,\mathbf{y}=\mathbf{e}$. We repeat this for $M$ independent draws of the white noise, resulting in a small ensemble $Y=\{\mathbf{y}^{(m)}\}^{M}_{m=1}$, with each field having the same covariance structure. We process these synthetic fields identically to the way we process the input ESM fields: we perform pixelwise standardization to ensure each pixel has a mean of 0 and a standard deviation of 1. The non-stationarity in our problem does not allow for more sophisticated normalization techniques [68], which either become computationally intractable, require a stationary structure, or are not applicable in this setting. An example spatial field and its associated parameter fields can be seen in Figure 6.
@@ C.2 Training
All networks are implemented in PyTorch [53] and trained on a single NVIDIA RTX A6000 GPU. The networks are trained using the AdamW optimizer [41] for 200 epochs with a step-wise learning rate decay and early stopping after 10 epochs without validation improvement. We use mean squared error (MSE) loss, computed on normalized parameter values within the training loop. This avoids loss imbalance caused by parameter scale differences while requiring less pre or post-processing from the user.
@@ Appendix D FURTHER EXPERIMENTS WITH SIMULATED DATA
RMSE $\downarrow$ MAE $\downarrow$ SSIM (to 1)($\uparrow\text${to }1) PSNR $\uparrow$ NRMSE $\downarrow$
@@
RMSE $\downarrow$ MAE $\downarrow$ SSIM (to 1)($\uparrow\text${to }1) PSNR $\uparrow$ NRMSE $\downarrow$
@@
RMSE $\downarrow$ MAE $\downarrow$ SSIM (to 1)($\uparrow\text${to }1) PSNR $\uparrow$ NRMSE $\downarrow$
@@ Appendix E FURTHER CLIMATE APPLICATION RESULTS
We use data from three climate models: the Max Planck Institute Earth System Model at Low Resolution (MPI-ESM 1.2LR), the Community Earth System Model (CESM1), and the European Community Earth System Model (EC-Earth3). We pass the standardized ensembles of temperature sensitivity anomalies into each of our estimation networks: STUN, UNet, ViT, CNN25, CNN17, and CNN9. Using the predicted parameters, we generate 1,000 synthetic fields from each model. Generation is performed using the LatticeKrig R package [50], which accommodates cylindrical geometry and ensures the correct treatment of spatial distances under the Mercator projection.
To evaluate how well each synthetic ensemble preserves the spatial correlation structure, we randomly select 50 anchor locations and compute the corresponding 50 rows from each simulated ensemble’s empirical correlation matrix. These are then compared to the 50 rows from the original ensemble’s correlation matrix using root mean squared error (RMSE). We systematically pair each I2I network with a local neural estimator of comparable rank (e.g., STUN vs. CNN25, UNet vs. CNN17), and compare their RMSEs across the same anchor locations. As the anchor locations are fixed for all pairings, we conduct paired t-tests between matched RMSEs to compare performance.
Our null hypothesis for each paired comparison is that the mean RMSE difference equals zero, $H_{0}\!:\,\mu_{d}=0$, where $d=\text{RMSE}_{\text{I2I}}-\text{RMSE}_{\text{CNN}}$. The one‑sided alternative, $H_{1}\!:\,\mu_{d}<0$, states that the I2I emulator attains a lower RMSE than its CNN counterpart. All tests are conducted with a significance level $\alpha=0.01$. Results are summarized in Table 5. For eight of the nine comparisons, the null hypothesis is rejected and the $99\%$ confidence interval is entirely negative, demonstrating that I2I networks tend to produce significantly more accurate correlation structure estimates than their corresponding local CNNs.
@@
Climate model Network Pair Paired t-test metrics
$\mu_{\text{I2I}}$ $\downarrow$ $\mu_{\text{local}}$ $\downarrow$ $\mu_{d}$ 99% CI $p_{\text{value}}$
MPI-ESM (192 $\times$ 96) STUN vs. CNN25 0.202 0.210 0.008-0.008 (-$\infty$, 0.004-0.004] $2.2\times 10^{-6}$
UNet vs. CNN17 0.203 0.214 0.011-0.011 (-$\infty$, 0.005-0.005] $1.0\times 10^{-5}$
ViT vs. CNN9 0.208 0.247 0.039-0.039 (-$\infty$, 0.028-0.028] $7.1\times 10^{-12}$
CESM1 (288 $\times$ 192) STUN vs. CNN25 0.229 0.243 0.013-0.013 (-$\infty$, 0.007-0.007] $1.1\times 10^{-6}$
UNet vs. CNN17 0.230 0.239 0.008-0.008 (-$\infty$, 0.004-0.004] $3.4\times 10^{-5}$
ViT vs. CNN9 0.228 0.244 0.016-0.016 (-$\infty$, 0.008-0.008] $5.9\times 10^{-6}$
EC-Earth3 (512 $\times$ 256) STUN vs. CNN25 0.271 0.285 0.014-0.014 (-$\infty$, 0.005-0.005] $1.6\times 10^{-4}$
UNet vs. CNN17 0.275 0.288 0.012-0.012 (-$\infty$, 0.004-0.004] $2.7\times 10^{-4}$
ViT vs. CNN9 0.283 0.304 0.021-0.021 (-$\infty$, $0.0003$] $1.1\times 10^{-2}$